ANALYSIS OF ROTATIONAL STRUCTURE IN THE HIGH-RESOLUTION INFRARED SPECTRA OF THE TRANS-HEXATRIENE-1,1-D_{2} AND -CIS-1-D_{1} SPECIES

\underline{NORMAN C. CRAIG}, HANNAH A. FUSON, and HENGFENG TIAN, Department of Chemistry and Biochemistry, Oberlin College, Oberlin, OH 44074; THOMAS A. BLAKE, Environmental Molecular Sciences Laboratory, Pacific Northwest National Laboratory, Richland, WA 99352.

Hexatriene-1,1-d_{2} with some admixture of the cis-1-d_{1} and trans-1-d_{1} species was synthesized by reaction of 2,4-pentadienal and (methyl-d_{3})-triphenylphosphonium iodide (Wittig reagent). The trans isomer was isolated by preparative gas chromatography, and the high-resolution (0.0015 cm^{−1}) infrared spectrum was recorded on a Bruker IFS 125HR instrument. The rotational structure in two C-type bands for the 1,1-d_{2} species was analyzed. For this species the bands at 902.043 and 721.864 cm^{−1} yielded composite ground state rotational constants of A_{0} = 0.801882(1), B_{0} = 0.041850(2), and C_{0} = 0.039804(1) cm^{−1}. For the cis-1-d_{1} species the C-type band at 803.018 cm^{−1} gave A_{0} = 0.809384(2), B_{0} = 0.043530(3), and C_{0} = 0.041321(2) cm^{−1}. By iodine-catalyzed isomerization, we have obtained some of the much less favored cis isomer and hope to obtain microwave spectra for its three deuterium-substituted species. The rotational constants reported here contribute to data needed for determining a semi-experimental structure for trans-hexatriene, which should show that the structural consequences of pi-electron delocalization increase with the chain length of polyenes.